# Building symmetries into data-driven manifold dynamics models for complex flows

### v1

@@ III.3 Time evolution of pattern with neural ODEs
In the previous two steps we first mapped our data to the fundamental chart, allowing us to represent the state with $\tilde{\omega}$, $\mathcal{I}$, and $\phi_{x}$. Then we reduced the dimension of $\tilde{\omega}$ by mapping it to the manifold coordinates $h$. Next, we need to find a method to evolve $h$, $\mathcal{I}$, and $\phi$ through time. To forecast $h(t)$ in time we use the neural ODE (NODE) framework[9, 33, 23]. We use a stabilized version proposed by Linot et. al. [33] where the dynamics on the manifold in the fundamental chart are described by the equation
@@
To train the NODE model we first gather $\omega(t)$ and $\omega(t+\tau)$. We then map to the fundamental chart in pairs. As an example, if a snapshot $\omega(t)$ lies in $\mathcal{I}=5$ we apply the corresponding discrete operations such that $h(t)$ falls in the interior of the fundamental chart $\mathcal{I}=0$. The same discrete operations, which are $\mathscr{S}\mathscr{R}$ (refer to Table 1), are applied to $\omega(t+\tau)$. This means that $h(t+\tau)$ does not need to fall in the interior. We select $\tau=0.5$ which is a small enough time such that the exterior region is covered by the autoencoder.
@@
Here we discuss the detailed operations for evolving trajectories in the fundamental chart with the NODE models. Fig. 10a summarizes the methodology for evolving in time. An initial condition in the fundamental space is first encoded and projected with $\hat{U}$ to obtain $h(t)$. Then the NODE maps a trajectory forward $T$ time units to yield $h_{\text{temp}}(t+T)$ (note the indicator does not change in this step). We select $T=\tau$. After this, we perform the appropriate symmetry operations, detailed below, to find $h_{r}(t+T)$ and update the new indicator $\mathcal{I}(t+T)$. This is repeated to continue evolving forward in time. In Fig. 10b we present the method for performing these symmetry operations.
There are two main steps in applying the symmetry operations: 1) determining $\mathcal{I}$ from $h_{\text{temp}}$, and 2) getting the values of $h_{r}(t+T)$ and $\mathcal{I}(t+T)$. Step 1 is simply a classification problem where we wish to find a function that classifies new values of $h_{\text{temp}}$ as either lying within or outside the fundamental domain. In Floryan & Graham [15] this was performed by identifying the nearest neighbor of the training data in the manifold coordinates and classifying $h_{\text{temp}}$ with the same label. However, any classification technique can work – for example, support vector machines (SVM) [34] also worked well at this task. In the case of our specific problem, we found the fastest method was to map $h_{\text{temp}}$ to the ambient space $\omega_{\text{temp}}$ at every step and calculate $\hat{\omega}_{R}(0,1)$, $\hat{\omega}_{I}(0,1)$, $\hat{\omega}_{I}(0,2)$ to verify if $h_{\text{temp}}$ lies in the fundamental space. Step 2 involves updating the manifold coordinates to get $h_{r}(t+T)$. If the indicator changes, symmetries are factored out with $\mathcal{I}(t+T)$ and the snapshot is encoded to get the new $h_{r}(t+T)$. If the indicator stays the same $h_{r}(t+T)=h_{\text{temp}}(t+T)$ and $\mathcal{I}(t+T)=\mathcal{I}(t)$.
Figure 10: (a) Time evolution of $h(t)$ with NODE and (b) symmetry operation check to map back to fundamental space if needed.
@@ III.4 Time evolution of phase with neural ODEs
The term $g_{\mathcal{I}}(\mathcal{I})$ equals 1 if an even number of discrete operations map $h$ back to the fundamental space, or -1 if an odd number of discrete operations map $h$ back to the fundamental space. To understand how the signs of Equation 14 change consider the effect of discrete symmetry operations on the phase calculation $\phi_{x}=\operatorname{atan}2\left\{\hat{\omega}_{I}(1,0),\hat{\omega}_{R}(1,0)\right\}$. Applying a discrete symmetry operation on a snapshot changes the phase variable to either
@@ IV Results
We present results for the chaotic case $\mathrm{Re}=14.4$, whose trajectories sample all eight charts introduced above. First we illustrate the effect of training data size on the autoencoder performance when considering the fundamental chart data, as opposed to using the original and phase-aligned data. These results are shown in Section IV.1.1. We then use IRMAE-WD to estimate the minimal dimension needed to represent the data. We summarize these results in Section IV.1.2. The time evolution model performance results are shown in Section IV.2. Here we confirm the equivariance of the model, (Section IV.2.1) then show the performance for short-time tracking, long-time statistics, and phase evolution (Sections IV.2.2, IV.2.3, and IV.2.4 respectively).
@@ IV.1 Dimension reduction with IRMAE-WD
To train our IRMAE-WD models we minimize $\mathcal{L}$ via stochastic gradient descent. Following Zeng et al. [12], we use the AdamW optimizer, which decouples weight decay from the adaptive gradient update and helps avoid the issue of weights with larger gradient amplitudes being regularized disproportionately, as observed in Adam [35]. All models were trained for a total of 1000 epochs and with a learning rate of $10^{-3}$. We consider $10^{5}$ snapshots of the original data ($\omega$), the phase-aligned data ($\omega_{l}$), and the data mapped to the fundamental space ($\tilde{\omega}$) separated by $\tau=0.5$ time units.
Before training the AEs, the mean is subtracted and the data is divided by its standard deviation. Three models are trained for each case of varying number of linear layers $L=0,4,6,8,10$ and weight decay values of $\lambda=10^{-4},10^{-6},10^{-8}$ for a total of 45 models for the $\omega$, $\omega_{l}$, and $\tilde{\omega}$ cases. For all of the networks discussed in this work, we use a $80/20$ split of the data for training and testing respectively. We select $d_{z}=40$, which is significantly higher than the dimension expected based on our previous work [10].
@@ IV.1.1 Effect of training data size on performance
A major benefit of mapping our data to the fundamental chart is that it results in eightfold denser sampling within that chart, as shown in Fig. 6. We also see that all the data is in a much smaller part of state space, and only that part of state space needs to be represented. As such, in this section, we test if this increased density allows us to use less data for training AEs. We factor out the symmetries of our system as discussed in Section III.1 and train IRMAE-WD models for the case of $L=4$ and $\lambda=10^{-4}$. In Fig. 11 we show the test MSE of the IRMAE-WD models as we vary the amount of data for three cases: the original data ($\omega$), the phase-aligned data ($\omega_{l}$), and the data mapped to the fundamental space ($\tilde{\omega}$). In all cases, we use the same $20,000$ test snapshots to calculate the MSE, train 3 models at each training data size, and reduce to a dimension of $d_{z}=40$. Notice that here we do not project to a lower dimension; that will be done below. The purpose here is not to map to the minimal dimension, but to study the effect of data size on training.
@@
As can be seen in Fig. 11, removing phase results in an order of magnitude improvement in the MSE over the original data, and mapping to the fundamental results in nearly another order of magnitude improvement in the MSE over the phase-aligned data. Thus, this drastic improvement in performance allows us to use far less data. For example, 800 snapshots in the fundamental space performs nearly as well as $80,000$ snapshots in the original space. Similarly, $8,000$ snapshots in the fundamental case outperforms $80,000$ snapshots in the phase-aligned case. Now that we have shown the effect of the number of snapshots on performance we use $80,000$ snapshots for the remaining studies.
@@
Figure 11: MSE vs number of data size used for training IRMAE-WD corresponding to a test data of 20,000 samples for $\text{Re}=14.4$. The parameters used are $L=4$ and $\lambda=10^{-4}$ and three trials are considered for each case. The data is reduced to a dimension of $d_{z}=40$.
IV.1.2 Effect on dimension estimates with varying linear layers and weight decay
@@
Here we study the effects on the dimension estimates when varying $L$ and $\lambda$ for the original, phase-aligned, and fundamental case. As discussed in Section III.2 we perform SVD on the covariance matrix of the encoded data $ZZ^{T}$ to obtain $U$ and truncate to obtain $h=\hat{U}^{T}z$. Zeng et al. [12] showed that structuring an autoencoder with linear layers and using weight decay causes the latent space to become low-dimensional through training. In Fig. 12, we show the evolution of these singular values through training. As the model trains for longer times, the trailing singular values tend towards zero. These can be truncated without any loss of information. For most cases, this drop is drastic ($\sim 10$ orders of magnitude) and a threshold can be defined to select how many singular values to keep (i.e. to select the number of dimensions).
@@
Figure 12: Evolution of singular values of the covariance matrix of the encoded test data $ZZ^{T}$ during training of an IRMAE-WD model with $L=4$ and $\lambda=10^{-4}$. Here the drop happens at $d_{h}=10$.
@@
Figure 13: MSE vs dimension $d_{h}$ given by the spectral gap of the singular values for the original (black), phase-aligned (red), and fundamental chart (blue) cases. Each case contains three trials of combinations of parameters $L=0,4,6,8,10$ and $\lambda=10^{-4},10^{-6},10^{-8}$.
@@
In Fig. 13 we plot the estimate of dimension for the original (black), phase-aligned (red), and fundamental chart (blue) data as we vary the number of linear layers $L$ and the weight decay parameter $\lambda$. For each case, there are two clear clusters – the cluster with higher dimension corresponds to $L=0$ and the other cluster has linear layers, $L>0$.This happens because in the absence of linear layers there is no mechanism to drive the rank to a minimal value. When we factor out symmetries the dimension estimates become less spread with ranges of $d_{h}=7-10$ for fundamental, $d_{h}=8-13$ for phase-aligned, and $d_{h}=11-18$ for original. This narrowing of the distribution likely happens due to the dense coverage of the state space in the fundamental chart, which better captures the shape of the manifold. The dimension estimate range from the fundamental chart is in good agreement with our previous observations of $d_{h}=9$ in [10]. In the following analysis we select a conservative estimate of the dimension which appears to be $d_{h}=10$ from the fundamental chart data.
@@ IV.2 Time evolution
To learn our NODE models we first train $g_{h}$ by minimizing Equation 13 ($\mathcal{L}_{\text{NODE}}$) and then we fix $g_{h}$ and train $g_{\phi}$ by minimizing Equation 14 ($\mathcal{L}_{\phi,\text{NODE}}$) via the Adam optimizer. We train for a total of 40000 epochs and a learning rate scheduler that drops from $10^{-3}$ to $10^{-4}$ at epoch number 13334 (1/3 into training) and from $10^{-4}$ to $10^{-5}$ at epoch number 26667 (2/3 into training). As previously discussed we use $\tau=0.5$ which ensures the trajectory spends some time steps in an overlap region as it moves from chart to chart, so we can learn the dynamics there.
@@ IV.2.1 Equivariance
where $i$ corresponds to the trajectory number, and $j$ to the initial chart (i.e. $j=0$ corresponds to the original initial condition of the $i$th trajectory). Fig. 14 presents this error through time for our symmetry charting method (Fundamental) and for models with only phase-alignment.
Figure 14: $TE(t)$ vs $t$ for Fundamental and Phase-aligned case where $TE(t)$ corresponds to the error calculation between predicted NODE trajectories, from a model with $d_{h}=10$, from initial conditions that are related by symmetry transformations.
As expected, the time integration from the model trained in the fundamental space satisfies equivariance perfectly with a $TE=0$. The phase-aligned curve does not satisfy equivariance, and we see that trajectories diverge fairly quickly.
A more severe consequence of not enforcing the discrete symmetries can be seen in trajectory predictions. Fig. 15 depicts $\|\omega(t)\|$ vs $t$ for the true data (shown in black) and the time integration of initial conditions starting in different charts (red – $\mathcal{I}=0$, blue – $\mathcal{I}=1$, and green – $\mathcal{I}=2$) for both fundamental (a) and phase-aligned cases (b). In all cases, the predicted trajectories diverge from the true trajectory after some time – this is expected for a chaotic system, as explore further in Section IV.2.2. However, the symmetry-charted predictions in Fig. 14(a) exhibit excellent short-time tracking and captures the bursting event that happens around $t=100$. At longer times, the prediction is not quantitatively accurate but still captures the alternation between quiescent and bursting intervals observed in the true data. By contrast, the predictions for the phase-aligned model, Fig. 14(b), deviate quickly from the true data, and furthermore do not even exhibit intermittency between quiescent and bursting dynamics – they stay in a bursting regime. Thus the models that do not account for the discrete symmetries do not capture the dynamics correctly, even at a qualitative level. These results reinforce the major advantage of properly accounting for symmetries, as the symmetry charting method does.
Figure 15: $\|\omega(t)\|$ vs time for initial conditions integrated with the (a) Fundamental and (b) Phase aligned NODE model for $d_{h}=10$. Colors correspond to initial conditions starting in different fundamental domains (red – $\mathcal{I}=0$, blue – $\mathcal{I}=1$, and green – $\mathcal{I}=2$). In (a), all the different-colored trajectories coincide.

### v2

@@ III.3 Time evolution of pattern with neural ODEs
In the previous two steps we first mapped our data to the fundamental chart, allowing us to represent the state with $\tilde{\omega}$, $\mathcal{I}$, and $\phi_{x}$. Then we reduced the dimension of $\tilde{\omega}$ by mapping it to the manifold coordinates $h$. Next, we need to find a method to evolve $h$, $\mathcal{I}$, and $\phi$ through time. To forecast $h(t)$ in time we use the neural ODE (NODE) framework[9, 37, 31], in particular a stabilized version proposed by Linot et. al. [37] where the dynamics on the manifold in the fundamental chart are described by the equation
@@
While in the present study we work with clean data from numerical simulations, the methods described here can be used in the presence of noise. In some cases, the dimension reduction step itself can serve as a noise filter. For example, linear dimension reduction with principal components analysis can be used as a preprocessing step to filter out noise [38]. Even in the presence of moderate noise, neural ODEs have been shown to perform well without changes to the training procedure [31, 39]. If the noise is severe, other steps can be taken, such as smoothing the data with a temporal filter or adding additional regularization to the neural ODE training [39, 40, 41].
Many studies have used variants of the Koopman operator formalism for data-driven time-evolution modeling [42, 43]. A widely used implementation of this approach has recently been shown to be equivalent to a NODE representation [44].
To train the NODE model we first gather $\omega(t)$ and $\omega(t+\tau)$. We then map to the fundamental chart in pairs. As an example, if a snapshot $\omega(t)$ lies in $\mathcal{I}=5$, we apply the corresponding discrete operation $\mathscr{R}\mathscr{S}$ such that $\tilde{\omega}(t)$ and thus $h(t)$ fall in the interior of the fundamental chart $\mathcal{I}=0$. The same discrete operation (refer to Table 1), is applied to $\omega(t+\tau)$. This means that $h(t+\tau)$ does not need to fall in the interior. We select $\tau=0.5$ which is a small enough time such that the exterior region is covered by the autoencoder.
@@
Here we discuss the detailed operations for evolving trajectories in the fundamental chart with the NODE models. Fig. 10a summarizes the methodology for evolving in time. An initial condition in the fundamental domain is first encoded and projected with $\hat{U}$ to obtain $h(t)$. Then the NODE maps a trajectory forward $T$ time units to yield $h_{\text{temp}}(t+T)$ (note the indicator does not change in this step). We select $T=\tau$. After this, we perform the appropriate symmetry operations, detailed below, to find $h_{r}(t+T)$ and update the new indicator $\mathcal{I}(t+T)$. This is repeated to continue evolving forward in time. In Fig. 10b we present the method for performing these symmetry operations.
There are two main steps in applying the symmetry operations: 1) determining $\mathcal{I}$ from $h_{\text{temp}}$, and 2) getting the values of $h_{r}(t+T)$ and $\mathcal{I}(t+T)$. Step 1 is simply a classification problem where we wish to find a function that classifies new values of $h_{\text{temp}}$ as either lying within or outside the fundamental domain. In Floryan & Graham [15] this was performed by identifying the nearest neighbor of the training data in the manifold coordinates and classifying $h_{\text{temp}}$ with the same label. However, any classification technique can work – for example, support vector machines (SVM) [45] also worked well at this task. In the case of our specific problem, we found the fastest method was to map $h_{\text{temp}}$ to the ambient space $\omega_{\text{temp}}$ at every step and calculate $\hat{\omega}_{R}(0,1)$, $\hat{\omega}_{I}(0,1)$, $\hat{\omega}_{I}(0,2)$ to verify if $h_{\text{temp}}$ lies in the fundamental domain. Step 2 involves updating the manifold coordinates to get $h_{r}(t+T)$. If the indicator changes, symmetries are factored out with $\mathcal{I}(t+T)$ and the snapshot is encoded to get the new $h_{r}(t+T)$. If the indicator stays the same $h_{r}(t+T)=h_{\text{temp}}(t+T)$ and $\mathcal{I}(t+T)=\mathcal{I}(t)$.
Figure 10: (a) Time evolution of $h(t)$ with NODE and (b) symmetry operation check to map back to the fundamental domain if needed.
@@ III.4 Time evolution of phase with neural ODEs
The term $g_{\mathcal{I}}(\mathcal{I})$ equals 1 if an even number of discrete operations map $h$ back to the fundamental domain, or -1 if an odd number of discrete operations map $h$ back to the fundamental domain. To understand how the signs of Equation 18 change consider the effect of discrete symmetry operations on the phase calculation $\phi_{x}=\operatorname{atan}2\left\{\hat{\omega}_{I}(1,0),\hat{\omega}_{R}(1,0)\right\}$. Applying a discrete symmetry operation on a snapshot changes the phase variable to either
@@ IV Results
We present results for the chaotic case $\mathrm{Re}=14.4$, whose trajectories sample all eight charts introduced above, as well as exhibiting a chaotically drifting phase and thus sampling all values of $\phi_{x}$. We focus on the comparative performance of approaches that do not exploit symmetry and those that do. First we illustrate the effect of training data size on the autoencoder performance when considering the fundamental chart data, as opposed to using the original and phase-aligned data. These results are shown in Section IV.1.1. We then use IRMAE-WD to estimate the minimal dimension needed to represent the data. We summarize these results in Section IV.1.2. The time evolution model performance results are shown in Section IV.2. Here we confirm the equivariance of the model, (Section IV.2.1) then show the performance for short-time tracking, long-time statistics, and phase evolution (Sections IV.2.2, IV.2.3, and IV.2.4 respectively).
@@ IV.1 Dimension reduction with IRMAE-WD
To train our IRMAE-WD models, we minimize the loss $\mathcal{L}$ in Eq. 14 via stochastic gradient descent. Following Zeng et al. [12], we use the AdamW optimizer, which decouples weight decay from the adaptive gradient update and helps avoid the issue of weights with larger gradient amplitudes being regularized disproportionately, as observed in Adam [46]. All models were trained for a total of 1000 epochs and with a learning rate of $10^{-3}$. The entire data set consists of a time series of $10^{5}$ snapshots at time intervals of $\tau=0.5$ time units. The first $80,000$ snapshots, or downsampled versions thereof as described below, are the training data, and the last $20,000$ are the test data. For cases where phase alignment is applied, the original training data snapshots $\omega$ are transformed as described above to $\omega_{l}$, and where the discrete symmetries are also factored out, they are mapped into the fundamental domain. This is then extended to form the fundamental chart; the data points in this chart are denoted $\tilde{\omega}.$ The operations for performing these transformations are described in Sec. III.1.
Before training the AEs, the mean is subtracted and the data is divided by its standard deviation. For each data set ($\omega$, $\omega_{l}$, and $\tilde{\omega}$ cases, with varying numbers of samples as described below), Three models are trained for each case of varying number of linear layers $L=0,4,6,8,10$ and weight decay values $\lambda=10^{-4},10^{-6},10^{-8}$. That is, for each data set we train 45 separate models. For all autoencoders, we select a bottleneck dimension $d_{z}=40$, which is significantly higher than the manifold dimension expected based on our previous work [10].
In Sec. IV.1.1, we examine the effect of symmetry reduction on the autoencoder error as we change the size of the test data set. In Sec. IV.1.2, we keep the data size constant and examine the effect of symmetry reduction on the estimated manifold dimension. In both cases, we will see that incorporating symmetry has a substantial benefit.
@@ IV.1.1 Effect of training data size on performance
A major benefit of mapping data to the fundamental chart is that it results in eightfold denser sampling within that chart, as shown by comparison of Figs. 5(a) and Fig. 6(a). We also see that all the data is in a much smaller part of state space, and only that part of state space needs to be represented. In this section, we test to what extent this increased density improves the autoencoder performance for a given set of training data. Using the full training data (80,000 snapshots at time intervals $\tau=0.5$), as well as reduced sets obtained by downsampling the training data set (increasing the time interval between snapshots), we train IRMAE-WD models with $L=4$ and $\lambda=10^{-4}$. This is done for the original data ($\omega$), the phase-aligned data ($\omega_{l}$), and the data mapped to the fundamental domain ($\tilde{\omega}$).
@@
In Fig. 11 we show the test MSE of the IRMAE-WD models, normalized based on the mean squared vorticity for the test data ($\omega_{\mathrm{test}}$), as we vary the amount of training data. As can be seen in Fig. 11, removing phase results in an order of magnitude improvement in the normalized MSE relative to the original data, and mapping to the fundamental results in nearly another order of magnitude improvement. This this drastic improvement in performance allows us to use far less data. For example, by comparing the blue diamonds to the black circles, we see that a model trained with only 800 snapshots in the fundamental domain performs nearly as well as one trained with $80,000$ snapshots in the original state space. Similarly, $8,000$ snapshots in the fundamental case outperforms $80,000$ snapshots in the phase-aligned case. We only show the performance for the IRMAE-WD architecture, but our approach only modifies the input data. This means it can be applied to any autoencoder architecture.
@@
Figure 11: Normalized autoencoder test MSE vs. training data size, $\text{Re}=14.4$. The parameters used are $L=4$ and $\lambda=10^{-4}$, and three trials are considered for each case.
IV.1.2 Effect of symmetry reduction and autoencoder parameters on manifold dimension estimate
@@
Here we study the effects on the dimension estimates when varying $L$ and $\lambda$ for the original, phase-aligned, and fundamental cases. The full training data set of 80,000 snapshots is used. As discussed in Section III.2 we perform SVD on the covariance matrix of the encoded data $ZZ^{T}$ to obtain $U$ and truncate to obtain $h=\hat{U}^{T}z$. Zeng et al. [12] showed that structuring an autoencoder with linear layers and using weight decay causes the latent space to become low-dimensional through training. In Fig. 12, we show an example of the evolution of these singular values through training, using data in the fundamental chart. As the model trains for longer times, the trailing singular values tend towards zero. These can be truncated without any loss of information. For most cases, this drop is drastic ($\sim 10$ orders of magnitude) and a threshold can be defined to select how many singular values to keep (i.e. to select the number of dimensions).
@@
Figure 12: Evolution of the spectrum of singular values $\sigma_{i}$ (ordered by magnitude) of the covariance matrix of the encoded test data $ZZ^{T}$ during training of an IRMAE-WD model on the symmetry-reduced data, with $L=4$ and $\lambda=10^{-4}$. Color indicates training epoch. Here the drop happens at $d_{h}=10$.
@@
Figure 13: Normalized MSE vs. dimension $d_{h}$ given by the spectral gap of the singular values for the original (black), phase-aligned (red), and fundamental chart (blue) cases. Each case contains three trials of combinations of parameters $L=0,4,6,8,10$ and $\lambda=10^{-4},10^{-6},10^{-8}$.
@@
In Fig. 13 we plot normalized MSE vs. estimate of dimension for the original (black), phase-aligned (red), and fundamental chart (blue) data as we vary the number of linear layers $L$ and the weight decay parameter $\lambda$. For each case, there are two clear clusters – the cluster with higher dimension corresponds to $L=0$ and the other cluster has linear layers, $L>0$. This happens because in the absence of linear layers there is no mechanism to drive the rank to a minimal value. When we factor out symmetries, the MSE decreases (as shown above), and the dimension estimates become less spread, with ranges of $d_{h}=7-10$ for fundamental, $d_{h}=8-13$ for phase-aligned, and $d_{h}=11-18$ for original. This narrowing of the distribution likely happens due to the dense coverage of the state space in the fundamental chart, enabling the autoencoder to better capture the shape of the manifold. The dimension estimate range from the fundamental chart is in good agreement with our previous observations of $d_{h}=9$ in [10]. In the following analysis, we select a conservative estimate of the dimension, which appears to be $d_{h}=10$ from the fundamental chart data.
@@ IV.2 Time evolution
To learn our NODE models we first train $g_{h}$ by minimizing Equation 17 for $\mathcal{L}_{\text{NODE}}$, and then we fix $g_{h}$ and train $g_{\phi}$ by minimizing Equation 18 ($\mathcal{L}_{\phi,\text{NODE}}$) via the Adam optimizer. We train for a total of 40000 epochs and a learning rate scheduler that drops from $10^{-3}$ to $10^{-4}$ at epoch number 13334 (1/3 into training) and from $10^{-4}$ to $10^{-5}$ at epoch number 26667 (2/3 into training). As previously discussed we use $\tau=0.5$ which ensures the trajectory spends some time steps in an overlap region as it moves from chart to chart, so we can learn the dynamics there.
@@ IV.2.1 Equivariance
where $i$ corresponds to the trajectory number, and $j$ to the initial chart (i.e. $j=0$ corresponds to the original initial condition of the $i$th trajectory). Fig. 14 presents this error through time for our symmetry charting method (blue) and for models with only phase-alignment (red).
Figure 14: $TE(t)$ vs. $t$ for Fundamental and Phase-aligned case where $TE(t)$ corresponds to the error calculation between predicted NODE trajectories, from a model with $d_{h}=10$, from initial conditions that are related by symmetry transformations.
As expected, the time integration from the model trained in the fundamental domain satisfies equivariance perfectly with a $TE=0$. The phase-aligned curve does not satisfy equivariance, and we see that trajectories diverge fairly quickly, leading to substantial error.
A more severe consequence of not enforcing the discrete symmetries can be seen in trajectory predictions. Fig. 15 depicts $\|\omega(t)\|$ vs. $t$ for the true data (shown in black) and the time integration of initial conditions starting in different charts (red – $\mathcal{I}=0$, blue – $\mathcal{I}=1$, and green – $\mathcal{I}=2$) for both fundamental (a) and phase-aligned cases (b). In all cases, the predicted trajectories diverge from the true trajectory after some time – this is expected for a chaotic system, as explore further in Section IV.2.2. However, the symmetry-charted predictions in Fig. 15(a) exhibit excellent short-time tracking and captures the bursting event that happens around $t=100$. At longer times, the prediction is not quantitatively accurate but still captures the alternation between quiescent and bursting intervals observed in the true data. By contrast, the predictions for the phase-aligned model, Fig. 15(b), deviate quickly from the true data, and furthermore do not even exhibit intermittency between quiescent and bursting dynamics – they stay in a bursting regime. Thus the models that do not account for the discrete symmetries do not capture the dynamics correctly, even at a qualitative level. These results reinforce the major advantage of properly accounting for symmetries, as the symmetry charting method does.
Figure 15: $\|\omega(t)\|$ vs. time for initial conditions integrated with the (a) Fundamental and (b) Phase aligned NODE model for $d_{h}=10$. Colors correspond to initial conditions starting in different fundamental domains (red – $\mathcal{I}=0$, blue – $\mathcal{I}=1$, and green – $\mathcal{I}=2$). In (a), all the different-colored trajectories coincide.